# REGRACE: A Robust and Efficient Graph-based Re-localization Algorithm using Consistency Evaluation

### v1

@@ Abstract
Loop closures are essential for correcting odometry drift and creating consistent maps, especially in the context of large-scale navigation. Current methods using dense point clouds for accurate place recognition do not scale well due to computationally expensive scan-to-scan comparisons. Alternative object-centric approaches are more efficient but often struggle with sensitivity to viewpoint variation. In this work, we introduce REGRACE, a novel approach that addresses these challenges of scalability and perspective difference in re-localization by using LiDAR-based submaps. We introduce rotation-invariant features for each labeled object and enhance them with neighborhood context through a graph neural network. To identify potential revisits, we employ a scalable bag-of-words approach, pooling one learned global feature per submap. Additionally, we define a revisit with geometrical consistency cues rather than embedding distance, allowing us to recognize far-away loop closures. Our evaluations demonstrate that REGRACE achieves similar results compared to state-of-the-art place recognition and registration baselines while being twice as fast. The code will be available upon acceptance.
@@ I INTRODUCTION
To provide a more geometrically accurate and lightweight representation, research works have used submaps as an input to SLAM [2] and re-localization pipelines [3, 4]. These dense partial maps are created by integrating LiDAR scans from multiple viewpoints into the same frame, which reduces the dependency on the perspective. Nonetheless, obtaining a compact yet informative representation of submaps remains challenging.
@@
This paper proposes REGRACE, a scalable place recognition and registration pipeline for submaps. We produce dense submaps by aggregating LiDAR scans into a single frame, expanding the sensor field of view and reducing object occlusion. We also propose context-aware object descriptors by representing each submap as a graph. Finally, we introduce a new approach to detecting loop closures, which uses geometric consistency to classify revisits rather than embedding distance, helping detect far-away revisits. Fig. 1 shows a typical result obtained with REGRACE. In summary, our contributions are:
@@ II-A Point cloud-based Place Recognition
Object-based localization addresses these issues by balancing reduced keypoint extraction with context-informed features [3, 15, 16]. SemSegMap [4] effectively identifies revisits using a CNN to encode each semantic segment while pooling a global descriptor using k-nearest neighbors. Several works [6, 7] improve the sensitivity to occlusion and viewpoint by representing the scene’s structure as edges in a graph, where each object instance is a node. However, instead of a global embedding per scan, these methods output similarity scores for pairs of scans, making them impractical for real-world robotics that use BoW querying.
To address the scalability and representation issues, REGRACE transforms each submap to a graph, where the vertices contain learned rotational-invariant descriptions for each object in the scene. We inform these local embeddings of their neighboring structure using a graph neural network. We also aggregate a global embedding using learned weights that balance the significance of each dimension of the local descriptors.
@@ III-B Overview
The REGRACE feature extractor pipeline is depicted in Fig. 2. We start by semantically classifying each collected LiDAR scan and fusing it into a larger submap. Next, we cluster the object instances and use a combination of convolutional and graph neural networks to extract and enhance a local descriptor for each object. Finally, we create a global descriptor by averaging all object features using a learning-based pooling. For place recognition, we retrieve the top-N closest global descriptors from a database of previously seen submaps and select the most geometrically consistent submap as a revisit candidate. To determine $\boldsymbol{T}_{\mathcal{S}_{\text{c}}\mathcal{S}_{j}}$ between the query $\mathcal{S}_{\text{c}}$ and the candidate match $\mathcal{S}_{j}\in\mathcal{S}$, we register the submaps using the local geometric-informed descriptors.
Figure 2: Overview of REGRACE feature extractor. First, it semantically labels each scan using Cylinder3D [24] and merges multiple scans into a submap using the poses retrieved by a SLAM pipeline. It then clusters the submap into object instances with DBSCAN [25], where each cluster is embedded using RIConv++ [26]. Finally, REGRACE polls a unified global descriptor from the node embeddings by leveraging a GNN. Feed-Forward Networks (FFN) connect the encoders.
@@ III-C Submap Generation
To ensure scalability to large maps, we discretize the submap $\mathcal{S}$ in ten-centimeter voxels. The semantic probability of a voxel is the average of the probabilities of all points inside it, smoothing out erroneous predictions for more accurate and consistent semantic information in each submap.
@@ III-D Local Features
We use the center of each object cloud $\boldsymbol{c}_{i}$ as keypoint coordinates, and learn object features $\boldsymbol{f}_{i}=\phi_{\mathcal{F}}(\mathcal{O}_{i})$ with $\phi_{\mathcal{F}}$ being a neural network. In REGRACE, $\phi_{\mathcal{F}}$ is based on RIConv++ [26], a rotational-invariant adaptation of PointNet++ [27]. Our network consists of a sparse convolutional encoder, which outputs object features $\boldsymbol{f}\in\mathbb{R}^{K\times 128}$, where $K$ is the number of objects in the submap. We sample $P=1,024$ farthest points for each cluster. If the point cloud $\mathcal{O}_{i}$ contains fewer points than $P$, we pad the available $p$ points with the $(P-p)$ farthest points in the cluster.
@@ III-E1 Graph Construction
$e_{ij}=\left\|\boldsymbol{c}_{i}-\boldsymbol{c}_{j}\right\|_{L2}/\alpha,$ (1)1( 1 )
@@ III-E2 Network Architecture
Being robust to viewpoint differences is essential to place recognition. Therefore, we enhance the local descriptors using an equivariant GNN (EGNN) [28], which is robust to the rotation and translation of the node positions. Our network outputs a graph $\tilde{\mathcal{G}}=\text{EGNN}(\mathcal{G})$ with updated nodes $\tilde{v}_{i}=\langle\tilde{\boldsymbol{c}}_{i},\tilde{\boldsymbol{f}}_{i}\rangle\in\tilde{\mathcal{V}}$, where $|\tilde{\mathcal{V}}|=\left|{\mathcal{V}}\right|=K$, $\tilde{\boldsymbol{c}}_{i}\in\mathbb{R}^{3}$ and ${\tilde{\boldsymbol{f}}}_{i}\in\mathbb{R}^{512}$. Note that $\tilde{\mathcal{G}}$ and ${\mathcal{G}}$ have the same number of nodes, but with higher-dimensional node embeddings and updated centroid positions. We pool a global descriptor per graph using generalized-mean pooling (GeM) [29]
$\boldsymbol{g}=\text{GeM}(\tilde{\mathcal{G}})=\left(\frac{1}{K}\sum_{i=1}^{K}(\tilde{\boldsymbol{f}}i)^{{\lambda}}\right)^{{{\lambda}^{-1}}},$ (2)2( 2 )
@@
$s(\boldsymbol{g}_{i},\boldsymbol{g}_{j})=\sigma\left(\text{ReLU}\left(\boldsymbol{g}_{i}^{T}\boldsymbol{\omega}^{[1:s]}\boldsymbol{g}_{j}+\boldsymbol{\alpha}\begin{bmatrix}\boldsymbol{g}_{i}\\ \boldsymbol{g}_{j}\end{bmatrix}+b\right)\right),$ (3)3( 3 )
@@ III-F Training Loss
$\mathcal{L}_{\text{triplet}}=\max\big{(}\|\boldsymbol{g}_{a}-\boldsymbol{g}_{p}\|_{L2}-\|\boldsymbol{g}_{a}-\boldsymbol{g}_{n}\|_{L2}+m\,,\,0\big{)},$ (4)4( 4 )
where $\boldsymbol{g}_{a}$, $\boldsymbol{g}_{p}$, $\boldsymbol{g}_{n}$ are the global embeddings of the anchor submap $\mathcal{S}_{a}$, positive $\mathcal{S}_{p}$ and negative examples $\mathcal{S}_{n}$, respectively. Like in LoGG3D-Net [8], we use as a positive example a submap $\mathcal{S}_{p}$ that is at most 3m from the anchor $\mathcal{S}_{a}$, whereas a negative submap $\mathcal{S}_{n}$ is at least 20m away.
@@
$\mathcal{L}_{\text{score}}=-\big{[}s_{ij}\log(l_{ij})+(1-s_{ij})\log(1-l_{ij})\big{]},$ (5)5( 5 )
@@
$\mathcal{L}=\mathcal{L}_{\text{triplet}}+\mathcal{L}_{\text{score}}.$ (6)6( 6 )
III-G Registration
REGRACE estimates the transformation between two submaps in a coarse-to-fine approach. To reduce the computation requirements, we initialize the transform via RANSAC [22] and then proceed to a more demanding ICP [23] step once an initial solution is obtained. For the RANSAC step, we realign only the object center coordinates $\boldsymbol{c}_{i}$ rather than the full point cloud. We match keypoints based on the corresponding node features from $\mathcal{G}$. A match happens if both nodes are mutually the best match between all the nodes of both graphs. Once the transform has been initialized using RANSAC $\boldsymbol{T}_{\mathcal{S}_{\text{c}}\mathcal{S}_{j}}^{\text{RANSAC}}$, we trigger an ICP optimization based on all points in the segmented clusters to retrieve a refined transformation $\boldsymbol{T}_{\mathcal{S}_{\text{c}}\mathcal{S}_{j}}^{\text{ICP}}$. To reduce the computational complexity of the ICP step, we register only the objects considered as inliers in the RANSAC step.
III-H Loop Closure Detection
@@
Previous place recognition pipelines classify scan pairs as loop closures if their embedding distance is below a predetermined threshold $\delta$. LoGG3D-Net classifies true revisits when the scans are positioned within 3m, while EgoNN [12] uses 5m and 20m. However, the relationship between global embeddings and actual distances is nonlinear, which makes a fixed threshold for identifying loop closure inefficient. To avoid triggering map optimization in areas that are not a loop closure, this threshold is often set lower than the pipeline’s capability of identifying hard revisits, increasing the number of false negatives. Moreover, submaps can have similar embeddings while still being farther apart than single LiDAR scans, leading to false positives.
@@
To reduce these false detections, we propose using a geometric verification criterion to classify loop closures. If the geometric consistency score $C(\mathcal{G}_{c},\mathcal{G}_{j})>\Delta$, the submap pair $(\mathcal{S}_{c},\mathcal{S}_{j})$ is considered a revisit. In addition to this new criterion, we re-rank the top 20 closest matches in the embedding space and select the one with the highest registration consistency score as the best-revisiting candidate. The consistency score $C(\mathcal{G}_{c},\mathcal{G}_{j})$ [17] is directly related to the registration success between the query submap $\mathcal{G}_{c}$ and submap $\mathcal{G}_{j}$ from the top-20 ($j\in\mathbb{Z}^{+},\forall j\leq 20$) closest embedding set
$C(\mathcal{G}_{c},\mathcal{G}_{j})=\sum_{\mathcal{I}}\text{max}\Bigg{(}1-\frac{(d_{1}-d_{2})^{2}}{(d_{t})^{2}},\,0\Bigg{)},$ (7)7( 7 )
@@
for $d_{1}=||\boldsymbol{c}_{i}-\boldsymbol{c}_{m}||_{L2}$ and $d_{2}=||\boldsymbol{c}_{j}-\boldsymbol{c}_{n}||_{L2}$, where $(v_{i},v_{j})\in\mathcal{G}_{c}$ and $(v_{m},v_{n})\in\mathcal{G}_{j}$ are a matching vertices. $\mathcal{I}$ is the set of inlier pairs of the RANSAC estimation, and $d_{t}=1$ is the maximum geometric error between the inlier pairs. For re-ranking we do not run the dense ICP step.
@@ IV-B Implementation Details
We trained REGRACE on a NVIDIA A40 GPU with Intel Xeon Gold 6254 @3.1GHz. The parameters used for DBSCAN are $\varepsilon=0.05$m and $\eta=800$. For KITTI-02, we used more flexible clustering parameters ($\varepsilon=0.1$m, $\eta=300$) as vegetation covers the background in most of the scans, making it difficult to gather large clusters. In the triplet loss $\mathcal{L_{\text{triplet}}}$, we used the margin $m=1$. The gradients are propagated end-to-end through the EGNN and the cluster feature extractor. We trained using Adam, with $10^{-4}$ learning rate over 100 epochs. On epochs 50 and 75, the learning rate decays by a factor of 10. In a final refinement step, we froze the RIConv++ network and trained only the EGNN over 50 epochs with a learning rate of $10^{-5}$.
@@ IV-C Metrics and Evaluation
We evaluate against LoGG3D-Net [8] following their 3m and 20m thresholds to classify true and false positives, respectively. We also compare to EgoNN [12], employing their scheme of 5m and 20m thresholds for true positives on KITTI-00 and 08. As an object-centric baseline, we compare to SGPR [6]. We report Recall@$1$, Recall@$5$, and $F1_{\text{max}}$. We iterate over values of $\delta$ to generate the precision and recall pairs for the $F1_{\text{max}}$.

### v2

@@ Abstract
Loop closures are essential for correcting odometry drift and creating consistent maps, especially in the context of large-scale navigation. Current methods using dense point clouds for accurate place recognition do not scale well due to computationally expensive scan-to-scan comparisons. Alternative object-centric approaches are more efficient but often struggle with sensitivity to viewpoint variation. In this work, we introduce REGRACE, a novel approach that addresses these challenges of scalability and perspective difference in re-localization by using LiDAR-based submaps. We introduce rotation-invariant features for each labeled object and enhance them with neighborhood context through a graph neural network. To identify potential revisits, we employ a scalable bag-of-words approach, pooling one learned global feature per submap. Additionally, we define a revisit with geometrical consistency cues rather than embedding distance, allowing us to recognize far-away loop closures. Our evaluations demonstrate that REGRACE achieves similar results compared to state-of-the-art place recognition and registration baselines while being twice as fast. Code and models are publicly available.
@@ I INTRODUCTION
To provide a more geometrically accurate and lightweight representation, research works have used submaps in SLAM [2] and re-localization [3, 4]. These dense partial maps are created by integrating LiDAR scans from multiple viewpoints into the same frame, which reduces the dependency on perspective. Nonetheless, obtaining a compact yet informative representation of submaps remains challenging.
@@
This paper proposes REGRACE, a scalable place recognition and registration pipeline for submaps. We produce dense submaps by aggregating LiDAR scans into a single frame, expanding the sensor field of view and reducing object occlusion. We also propose context-aware object descriptors by representing each submap as a graph , and introduce a new approach to detecting loop closures, which uses geometric consistency to classify revisits rather than embedding distance, helping detect far-away revisits. Fig. 1 shows a typical result obtained with REGRACE. In summary:
@@ II-A Point cloud-based Place Recognition
Object-based localization addresses these issues by balancing reduced keypoint extraction with context-informed features [3, 15, 16]. SemSegMap [4] effectively identifies revisits using a CNN to encode each semantic segment while pooling a global descriptor using k-nearest neighbors. Several works improve the sensitivity to occlusion and viewpoint by representing the scene’s structure in bird’s eye view [17] or as edges in a graph [6, 7], where each object instance is a node. However, instead of a global embedding per scan, these methods output similarity scores for pairs of scans, making them impractical for real-world robotics that use BoW querying.
To address the scalability and representation issues, we transform each submap to a graph, where the vertices contain learned rotational-invariant descriptions for each object in the scene. We inform these local embeddings of their neighboring structure using a graph neural network. We also aggregate a global embedding using learned weights that balance the significance of each dimension of the local descriptors.
@@ III-B Overview
The REGRACE feature extractor is depicted in Fig. 2. We start by semantically classifying each LiDAR scan and fusing it into a larger submap. Next, we cluster the objects and use convolutional and graph neural networks to extract a local descriptor for each instance. Finally, we create a global descriptor by averaging all object features using a learning-based pooling. For place recognition, we retrieve the top-N closest global descriptors from a database of previously seen submaps and select the most geometrically consistent submap as a revisit candidate. To determine $\boldsymbol{T}_{\mathcal{S}_{\text{c}}\mathcal{S}_{j}}$ between the query $\mathcal{S}_{\text{c}}$ and the candidate match $\mathcal{S}_{j}\in\mathcal{S}$, we register the submaps using the local geometric-informed descriptors.
Figure 2: Overview of REGRACE feature extractor. It semantically labels each scan using Cylinder3D [25] and merges multiple scans into a submap using the poses retrieved by a SLAM pipeline. It then clusters the submap into object instances with DBSCAN [26], where each cluster is embedded using RIConv++ [27]. Finally, REGRACE polls a unified global descriptor from the node embeddings. Feed-Forward Networks (FFN) connect the encoders.
@@ III-C Submap Generation
To ensure scalability to large maps, we discretize the submap $\mathcal{S}$ in ten-centimeter voxels. The semantic probability of a voxel is the average of the probabilities of all points inside it, smoothing out erroneous predictions.
@@ III-D Local Features
[…]
@@ III-E1 Graph Construction
$e_{ij}=\left\|\boldsymbol{c}_{i}-\boldsymbol{c}_{j}\right\|_{L2}/\alpha,$ (1)( 1 )
@@ III-E2 Network Architecture
Being robust to viewpoint differences is essential to place recognition. We enhance the local descriptors using an equivariant GNN (EGNN) [29], which is robust to rotation and translation of the node positions. Our network outputs a graph $\tilde{\mathcal{G}}=\text{EGNN}(\mathcal{G})$ with nodes $\tilde{v}_{i}=\langle\tilde{\boldsymbol{c}}_{i},\tilde{\boldsymbol{f}}_{i}\rangle\in\tilde{\mathcal{V}}$, where $|\tilde{\mathcal{V}}|=\left|{\mathcal{V}}\right|=K$, $\tilde{\boldsymbol{c}}_{i}\in\mathbb{R}^{3}$ and ${\tilde{\boldsymbol{f}}}_{i}\in\mathbb{R}^{512}$. Note that $\tilde{\mathcal{G}}$ and ${\mathcal{G}}$ have the same number of nodes, but with higher-dimensional node embeddings and updated centroid positions. We pool a global descriptor per graph using generalized-mean pooling (GeM) [30]
$\boldsymbol{g}=\text{GeM}(\tilde{\mathcal{G}})=\left(\frac{1}{K}\sum_{i=1}^{K}(\tilde{\boldsymbol{f}}i)^{{\lambda}}\right)^{{{\lambda}^{-1}}},$ (2)( 2 )
@@
$s(\boldsymbol{g}_{i},\boldsymbol{g}_{j})=\sigma\left(\text{ReLU}\left(\boldsymbol{g}_{i}^{T}\boldsymbol{\omega}^{[1:s]}\boldsymbol{g}_{j}+\boldsymbol{\alpha}\begin{bmatrix}\boldsymbol{g}_{i}\\ \boldsymbol{g}_{j}\end{bmatrix}+b\right)\right),$ (3)( 3 )
@@ III-F Training Loss
$\mathcal{L}_{\text{triplet}}=\max\big{(}\|\boldsymbol{g}_{a}-\boldsymbol{g}_{p}\|_{L2}-\|\boldsymbol{g}_{a}-\boldsymbol{g}_{n}\|_{L2}+m\,,\,0\big{)},$ (4)( 4 )
where $\boldsymbol{g}_{a}$, $\boldsymbol{g}_{p}$, $\boldsymbol{g}_{n}$ are the global embeddings of the anchor $\mathcal{S}_{a}$, positive $\mathcal{S}_{p}$ and negative $\mathcal{S}_{n}$ submaps, respectively. Like in LoGG3D-Net [8], we consider $\mathcal{S}_{p}$ at most 3m from $\mathcal{S}_{a}$, whereas $\mathcal{S}_{n}$ is at least 20m away from $\mathcal{S}_{a}$.
@@
$\mathcal{L}_{\text{score}}=-\big{[}s_{ij}\log(l_{ij})+(1-s_{ij})\log(1-l_{ij})\big{]},$ (5)( 5 )
@@
$\mathcal{L}=\mathcal{L}_{\text{triplet}}+\mathcal{L}_{\text{score}}.$ (6)( 6 )
III-G Loop Closure Detection
@@
Previous place recognition pipelines classify scan pairs as loop closures if their embedding euclidean distance $D$ is below a predetermined threshold $\delta$. LoGG3D-Net classifies true revisits when the scans are positioned within 3m, while EgoNN [12] uses 5m and 20m. However, the relationship between global embeddings and actual distances is nonlinear, which makes a fixed threshold for identifying loop closure inefficient. To avoid triggering map optimization in areas that are not a loop closure, this threshold $\delta$ is often set lower than the pipeline’s capability of identifying hard revisits, increasing the number of false negatives. Moreover, submaps can have similar embeddings while still being farther apart than single LiDAR scans, which leads to false positives.
@@
To reduce these false detections, we propose using a geometric criterion to classify loop closures. If the geometric consistency score $C(\mathcal{G}_{i},\mathcal{G}_{j})>\varepsilon$, the submap pair $(\mathcal{S}_{i},\mathcal{S}_{j})$ is considered a revisit. Note that the minimum consistency $\varepsilon$ between two graphs differs from the maximum distance between their global descriptors in the embedding space $\delta$. For each new submap, we re-rank the top 20 closest previous submaps based on embedding similarity $D$, and then select the one with the highest geometric consistency $C$ as the best revisiting candidate. The consistency score $C(\mathcal{G}_{i},\mathcal{G}_{j})$ [18] is directly related to the registration success between the query submap $\mathcal{G}_{i}$ and a submap $\mathcal{G}_{j}$ in the top 20 closest embeddings as in
$C(\mathcal{G}_{i},\mathcal{G}_{j})=\sum_{\mathcal{I}}\text{max}\Bigg{(}1-\frac{(d_{1}-d_{2})^{2}}{(d_{t})^{2}},\,0\Bigg{)},$ (7)( 7 )
@@
for $d_{1}=||\boldsymbol{c}_{i}-\boldsymbol{c}_{j}||_{L2}$ and $d_{2}=||\boldsymbol{c}_{i^{\prime}}-\boldsymbol{c}_{j^{\prime}}||_{L2}$, where $(v_{i},v_{i}^{\prime})\in\mathcal{G}_{i}$ and $(v_{j},v_{j}^{\prime})\in\mathcal{G}_{j}$ are a matching pair of vertices from RANSAC [23], $\mathcal{I}$ is the set of inlier pairs of this RANSAC estimation, and $d_{t}$ is the maximum geometric error between the inlier pairs. For re-ranking we do not run the dense ICP step. We assume $d_{t}=1$. The RANSAC realignment is further described in the following section.
III-H Registration
REGRACE estimates the transformation between two submaps in a coarse-to-fine approach. To reduce the computation requirements, we initialize the transformation via RANSAC and then proceed to a more demanding ICP [24] step once an initial solution is obtained. For the RANSAC step, we realign the object center coordinates $\boldsymbol{c}_{i}$ rather than the full point cloud. We match keypoints based on the corresponding node features from $\mathcal{G}_{i}$ and $\mathcal{G}_{j}$. A match happens if both nodes are mutually the best match between all the nodes of both graphs. Once the transform $\boldsymbol{T}_{\mathcal{S}_{\text{c}}\mathcal{S}_{j}}^{\text{RANSAC}}$ has been initialized using RANSAC, we trigger an ICP optimization based on all points in the segmented clusters to retrieve a refined transformation $\boldsymbol{T}_{\mathcal{S}_{\text{c}}\mathcal{S}_{j}}^{\text{ICP}}$. To reduce the computational complexity of the ICP step, we register only the objects considered as inliers in the RANSAC step.
@@ IV-B Implementation Details
We trained REGRACE on a NVIDIA A40 GPU with Intel Xeon Gold 6254 @3.1GHz. The parameters used for DBSCAN were $\varepsilon=0.05$m and $\eta=800$. For KITTI-02, we used more flexible parameters ($\varepsilon=0.1$m, $\eta=300$) as vegetation covers the background in most of the scans, making it difficult to gather large clusters. In the triplet loss $\mathcal{L_{\text{triplet}}}$, we used the margin $m=1$. We trained using Adam, with $10^{-4}$ learning rate over 100 epochs. On epochs 50 and 75, the learning rate decays by a factor of 10. In a final refinement step, we froze the RIConv++ network and trained only the EGNN over 50 epochs with a learning rate of $10^{-5}$.
@@ IV-C Metrics and Evaluation
We evaluate against LoGG3D-Net [8] following their 3m and 20m thresholds to classify true and false positives, respectively. We also compare to EgoNN [12], employing their scheme of 5m and 20m thresholds for true positives on KITTI-00 and 08. The lower the true positive threshold, the closer two submaps must be to be considered a match. By increasing this threshold, we can assess how far in advance the algorithm can identify a potential loop closure. As an object-centric baseline, we compare to SGPR [6]. We report Recall@111, Recall@555, and $F1_{\text{max}}$. We iterate over values of $\delta$ to generate the precision and recall pairs for the $F1_{\text{max}}$.